### v2

# MMedAgent-RL: Optimizing Multi-Agent Collaboration for Multimodal Medical Reasoning

Abstract
Medical Large Vision-Language Models (Med-LVLMs) have shown strong potential in multimodal diagnostic tasks. However, existing single-agent models struggle to generalize across diverse medical specialties, limiting their performance. Recent efforts introduce multi-agent collaboration frameworks inspired by clinical workflows, where general practitioners (GPs) and specialists interact in a fixed sequence. Despite improvements, these static pipelines lack flexibility and adaptability in reasoning. To address this, we propose MMedAgent-RL, a reinforcement learning (RL)-based multi-agent framework that enables dynamic, optimized collaboration among medical agents. Specifically, we train two GP agents based on Qwen2.5-VL via RL: the triage doctor learns to assign patients to appropriate specialties, while the attending physician integrates the judgments from multi-specialists and its own knowledge to make final decisions. To address the inconsistency in specialist outputs, we introduce a curriculum learning (CL)-guided RL strategy that progressively teaches the attending physician to balance between imitating specialists and correcting their mistakes. Experiments on five medical VQA benchmarks demonstrate that MMedAgent-RL not only outperforms both open-source and proprietary Med-LVLMs, but also exhibits human-like reasoning patterns. Notably, it achieves an average performance gain of 20.7% over supervised fine-tuning baselines.
1 Introduction
Large Vision-Language Models (LVLMs) are becoming increasingly proficient in visual understanding and reasoning [30, 31, 64, 2, 8]. This advancement is also making a significant impact in the biomedical domain, where Medical Large Vision-Language Models (Med-LVLMs) have demonstrated great potential in enabling intelligent diagnostic applications [24, 36, 38, 56, 6]. However, as shown in Figure 1 (a) left, although a single Med-LVLM can be trained with a large amount of data and show promise results to some extent, it is challenging for a single model to handle diagnostic expertise from different subfields (e.g., radiology, pathology, etc.).
Therefore, some recent works propose using multi-agent collaboration [55, 26, 16, 14] to solve medical tasks [27, 20, 48], where different models act as specialists or general practitioners, collaborating and discussing to arrive at a final answer, improving overall performance compared to a single agent. These works follow the steps of simulating a hospital visit process and adopt a General Practitioner (GP) $\rightarrow$ Specialist $\rightarrow$ GP workflow. First, the general practitioner (i.e., the triage doctor) classifies the patient based on the consultation questions and images and selects the appropriate department from several predefined specialties. Then, specialist doctors from the relevant departments provide their diagnoses. Finally, the general practitioner (i.e., attending physician) makes the final decision based on the images, consultation questions, and the diagnostic results from multiple specialists. However, as illustrated in Figure 1 (a) middle, such workflows are inherently static. Such interaction pattern between agents is fixed and predetermined, which limits the system’s capacity for flexible, optimized reasoning across multiple modalities.
Figure 1: Comparison of Med-Agent paradigms: single-agent $\rightarrow$ static workflows $\rightarrow$ dynamic collaboration. (a) Motivation: Single-agent models struggle with domain specialization, and prior multi-agent systems rely on fixed workflows, limiting adaptability. We propose a trainable reasoning-enhanced multi-agent system via RL. (b) Performance: Our method is highly competitive across multiple benchmarks.
To address this challenge, motivated by the success of Reinforcement Learning (RL) [47]-driven reasoning [19, 13, 50], as shown in Figure 1 (a) right, we perform a stage-wise training on two GPs based on Qwen2.5-VL [3] via reinforcement learning, namely the triage doctor and the attending physician. Specifically, first, for the first GP, i.e., triage doctor, we utilize the image modality information provided by the dataset itself to reinforce the triage doctor, such as pathology slides $\rightarrow$ Pathologist, ensuring that the triage doctor can accurately assign patients to the appropriate department. Then, we use powerful proprietary models like GPT-4o [39] to play the role of the specialist doctors and generate initial judgments. Finally, the second GP, i.e., attending physician, integrates domain knowledge from multiple specialists and their own judgment to make the final decision. Here, during the process of the general practitioner integrating specialist doctor information, while specialist doctors provide valuable domain knowledge, their judgments are not always perfectly accurate. These inconsistencies in specialist performance can introduce noise into model training, preventing the model from simply memorizing or blindly replicating their outputs. Instead, the model must learn to generalize beyond potentially flawed expert judgments. To address this, inspired by Curriculum Learning (CL) [4, 42, 10], which enables models to be trained progressively on increasingly difficult tasks, we implement a reinforcement learning approach based on CL, aiming to help the model gradually learn to leverage the knowledge of specialist doctors. This stepwise learning facilitates better knowledge integration and application by first teaching the model to simply imitate the answers of specialist doctors, and then gradually guiding it to reason out the correct answers when the specialists make mistakes. Specifically, we use the accuracy of specialist results as a flag to classify the training data by difficulty: specialist results that are completely correct are labeled as easy, partially correct as medium, and completely incorrect as hard. In this way, we design a three-stage curriculum reinforcement learning process for optimizing the attending physician to handle diverse specialist results, including how to accurately reference specialist knowledge and when to rely on its own understanding to solve problems.
The primary contribution of this paper is MMedAgent-RL, an RL-driven framework optimized for multi-agent collaboration in improving medical reasoning. Empirical results on five medical multimodal datasets, shows that the model performs exceptionally well not only on in-domain datasets but also on out-of-domain datasets, outperforming a series of both open-source and proprietary LVLMs, exceeding SFT method [40] by 20.7%. In addition to the quantitative performance, we also observed that the model’s reasoning process resembles human doctors’ thought processes: it first defines the disease mentioned in the question, carefully analyzes the image, and then determines if it can fully match the disease definition.
Figure 2: Overview of MMedAgent-RL, a RL-driven multi-agent framework designed to enhance the multimodal medical reasoning. It simulates the clinical loop of General Practitioner (GP) $\rightarrow$ Specialists $\rightarrow$ GP. MMedAgent-RL uses GRPO [13] to optimize the triage doctor (the first GP) in order to improve triage accuracy. Then, powerful proprietary LVLMs are used as the specialist doctors for the assigned department. Finally, curriculum learning [4, 42] and RL are combined to progressively train the attending physician (the second GP), who integrates the diverse opinions of specialists and makes robust decisions under varying levels of expert reliability.
2 Preliminaries
In this section, we will provide a brief overview of LVLMs, multi-agent collaboration and GRPO.
Large Vision Language Models. LVLMs enhance LLMs by integrating visual input $x_{v}$ with textual input $x_{t}$, forming a joint input $x=(x_{v},x_{t})$. They autoregressively predict the next token’s distribution to generate a textual response $y$.
Multi-Agent Collaboration. To support complex workflows, multi-agent frameworks coordinate specialized agents. Our setting simulates a hospital visit: GP $\rightarrow$ Specialist(s) $\rightarrow$ GP. Each agent $a_{i}\in\mathcal{A}$ follows policy $\pi_{\theta_{i}}(y\mid x)$, with multimodal input $x=(x_{v},x_{t})$, where $x_{v}$ is an image, $x_{t}$ is a text instruction, and $y$ is the output. GP agent: $a_{\text{GP}}$; specialists: $\{a_{\text{SP}}^{(1)},\ldots,a_{\text{SP}}^{(K)}\}$. The workflow proceeds as follows: 1) Triage: $a_{\text{GP}}^{\text{triage}}$ selects department via $d=\arg\max_{k}\pi_{\theta_{\text{GP}}^{\text{triage}}}(k\mid x)$. 2) Specialist: $a_{\text{SP}}^{(d)}$ produces response $y_{d}\sim\pi_{\theta_{\text{SP}}^{(d)}}(y\mid x)$. 3) Aggregation: $a_{\text{GP}}^{\text{attend}}$ outputs $y_{\text{final}}\sim\pi_{\theta_{\text{GP}}^{\text{attend}}}(y\mid x,y_{d})$.
Group Relative Policy Optimization (GRPO). Group Relative Policy Optimization (GRPO) [13] is a reinforcement learning method that avoids training a critic by using intra-group relative rewards to optimize the policy. For each query $x$, the model samples $G$ responses $\{y^{(1)},\ldots,y^{(G)}\}$, which are scored to get rewards $\{R_{1},\ldots,R_{G}\}$. GRPO computes normalized advantages and updates the policy with a PPO-style clipped objective [44]:
$\mathcal{J}_{\text{GRPO}}(\theta)=\mathbb{E}_{x,\{y_{i}\}}\Bigg{[}\frac{1}{G}\sum_{i=1}^{G}\Bigg{(}\min\Big{(}r_{i}A_{i},\,\text{clip}(r_{i},1-\epsilon,1+\epsilon)A_{i}\Big{)}-\beta\,\mathbb{D}_{\text{KL}}(\pi_{\theta}\|\pi_{\text{ref}})\Bigg{)}\Bigg{]},r_{i}=\frac{\pi_{\theta}(y_{i}\mid x)}{\pi_{\text{old}}(y_{i}\mid x)},$ (1)
where $A_{i}=\frac{R_{i}-\text{mean}(\{R_{j}\}_{j=0}^{G})}{\text{std}(\{R_{j}\}_{j=0}^{G})}$, $\epsilon$, $\beta$ are hyperparams, and $\pi_{\text{old}}$ is a old policy model. GRPO enables scalable policy learning using only relative rewards, without a critic.
Fine-tuning with GRPO. We fine-tune our model $\pi_{\theta}$ using GRPO to enhance the reasoning process. Each training sample $x\in\mathcal{D}$ consists of an image $x_{v}$, a textual prompt $x_{t}$ composed of a user query and system message, and a reference output format containing <think> and <answer> tags. Given input $x=(x_{v},x_{t})$, the model samples a group of candidate outputs $\{o^{(1)},\ldots,o^{(G)}\}\sim\pi_{\theta}(y\mid x)$, where each $o^{(i)}$ includes both reasoning trace and final answer. A reward function scores each output based on correctness of the final answer $r_{\text{accuracy}}$ and formatting compliance with XML-style tags $r_{\text{format}}$. Relative scores $\{R_{i}\}$ are computed and normalized to obtain advantages $\{A_{i}\}$, which are used to update $\pi_{\theta}$ following the GRPO objective (Eq. (1)).
3 Methodology
In this section, as illustrated in Figure 2, we will present MMedAgent-RL, a novel RL-driven multi-agent framework for multimodal medical reasoning by emulating a structured clinical workflow. Our approach begins with the first General Practitioner (GP) leveraging the input information to intelligently select the most appropriate medical department for further consultation. To optimize the accuracy of this initial triage decision, we employ GRPO [13] to train and refine the triage doctor’s capabilities. Subsequently, the case is referred to a panel of specialist doctors, each represented by a powerful proprietary LVLM specialized in the identified department. These LVLMs analyze the provided multimodal data and generate expert opinions relevant to their respective fields. Finally, the process culminates with the second GP, acting as the attending physician, who integrates the diverse insights from the specialist LVLMs along with their own medical knowledge to arrive at a comprehensive and robust final judgment. We will delve into the specifics of each stage as follows:
3.1 Initial Departmental Triage
In real-world medical treatment processes, the first critical step when a patient visits a hospital is triage [1], which involves assigning the patient to the appropriate department based on their symptoms. Previous works simulated hospital visits using LLMs or LVLMs, with Agent Hospital [27] employing an LLM for triage, and MDAgents [20] enhancing this by dynamically selecting collaboration structures and using GPT-4o for department assignment. However, these methods rely on a rigid and predefined approach for department assignment as they cannot update the model based on new data.
To address this challenge, in our approach, the first step is to optimize the general practitioner $a_{\text{GP}}^{\text{triage}}$ who acts as the triage doctor (i.e., policy $\pi_{\theta_{\text{GP}}^{\text{triage}}}$), using Qwen2.5-VL [3] as the base model. Here, we use the image modality information provided by the dataset itself as ground truth labels $y^{*}$ to train the triage model. For example, pathology slides $\rightarrow$ pathologist (e.g., PathVQA contains pathology slide images and is thus assigned to pathologists), chest X-ray images $\rightarrow$ radiologist (e.g., VQA-RAD includes various types of radiological images and is therefore assigned to radiologists), ensuring that the triage model can accurately assign patients to the appropriate medical specialty.
Specifically, when prompting the triage doctor, we provide $k$ candidate specialties. In our setup, $k$ is set to 7, including Pathologist, Radiologist, Surgeon, Oncologist, Endocrinologist, Ophthalmologist, and Dermatologist, which broadly cover the main departments involved in the data. Our aim is not only to improve triage accuracy, but also to strengthen the model’s reasoning process, helping explain why a particular triage recommendation is made. Thus, we use GRPO [13] as the base RL algorithm. At this stage, the reward function adopts a rule-based format with rewards $R_{\text{format}}\in\{0,0.5\}$ and accuracy rewards $R_{\text{accuracy}}\in\{0,1\}$. A reward is given when the output format meets the required criteria and the chosen specialty is correct. This stage can optimize the triage doctor’s performance, improving their ability to select the appropriate specialty $\small d=\arg\max_{k}\pi_{\theta_{\text{GP}}^{\text{triage}}}(k\mid x)$, and lays a foundation for subsequently acting as the corresponding specialist.
3.2 Role-Playing Specialists Offer Valuable Insights
After obtaining the department from the triage doctor, following previous work using LLMs or LVLMs for medical discussions [27, 48, 23, 49], we utilize several powerful models as specialist doctors $\small a_{\text{SP}}^{(d)}$ to provide relatively accurate preliminary judgments. This facilitates subsequent reference by the attending physician. In our setup, we use responses from $e$ specialists as references for each sample. We only require the specialist doctors to independently provide expert opinions $\small y_{d}\sim\pi_{\theta_{\text{SP}}^{(d)}}(y\mid x)$ within their specialty, without engaging in complex interactions. This ensures system efficiency and avoids majority voting that could overshadow minority opinions, leaving the final decision to the attending physician.
3.3 Evolving Decisions by Attending Physician via Ongoing Collaboration
Figure 3: (a) The average performance of both general practitioners (GP) and specialists is below 70%, which suggests a misalignment issue in multi-agent collaboration. Over-reliance on specialists’ opinions or unilateral decisions by GPs can both lead to suboptimal outcomes. (b) Since specialists perform inconsistently across different cases, this poses a challenge for GPs when making decisions. Using all data for reinforcement fine-tuning can easily trap the model in a locally suboptimal solution (left). In contrast, our C-MARL approach enables the model to progressively accomplish sub-goals in a three-stage process and ultimately reach a globally optimal solution (right).
After getting the responses from the specialists, we then integrate their knowledge into the final general practitioner designed to support the final diagnostic decision. The final decision-making agent, namely the attending physician, plays the most crucial role throughout the diagnostic process, as they must synthesize diverse expert opinions and draw upon their own clinical expertise to arrive at a final judgment. This poses significant challenges for the attending physician, as the specialists’ conclusions are not always fully reliable. For example, as shown in Figure 3 (a), although specialists generally outperform general practitioners, their overall accuracy is still below 70%. Therefore, over-reliance on specialist input can lead to suboptimal outcomes. Secondly, different specialists may offer conflicting interpretations of the same case, creating misalignment issues. If the model is unable to reconcile its internal reasoning with external expert input, it risks compounding errors. For instance, while majority voting may help mitigate the influence of less competent specialists, it can also suppress minority views—including potentially the only correct one. As illustrated in Figure 3 (a), such multi-agent collaboration (GPT-4o $\rightarrow$ Qwen2.5-VL or GPT-4o) can yield adverse effects when the model is not properly aligned with the nature and limitations of expert knowledge.
To address these challenges, as shown in Figure 3 (b), we draw inspiration from curriculum learning [4, 42, 37], which emphasizes the importance of organizing learning experiences in a meaningful progression, i.e., from easier to harder tasks. Motivated by this principle, we propose the Curriculum-based Reinforcement Learning tailored for Multi-Agent Collaboration (C-MARL), which gradually increases task difficulty during training. This stepwise learning approach promotes better knowledge integration and application by first training the model to mimic specialist doctors’ answers, and then gradually encouraging it to reason independently and correct errors when specialists make mistakes.
Specifically, unlike previous curriculum learning approaches that define difficulty based on problem formulation or data domains, we categorize tasks based on the accuracy of specialists’ diagnoses $\small y_{d}\sim\pi_{\theta_{\text{SP}}^{(d)}}(y_{d}\mid x)$, denoted by $s=\text{Acc}(y_{d},y^{*})$. The dataset is divided into three levels: fully correct specialist results ($s=1$) are labeled as easy, partially correct results ($10<s<$ 1) are labeled as medium, and completely incorrect results ($s=0$) are labeled as hard. The datasets corresponding to the three levels are denoted as $\mathcal{D}_{\text{easy}},\mathcal{D}_{\text{medium}},\mathcal{D}_{\text{hard}}$, respectively, and $\mathcal{D}=\mathcal{D}_{\text{easy}}\cup\mathcal{D}_{\text{medium}}\cup\mathcal{D}_{\text{hard}}$. Based on these data of three categories, as shown in Algorithm 1, we design a three-stage curriculum reinforcement learning process to optimize the attending physician’s ability to handle different types of specialist knowledge, such as when to accurately leverage specialist knowledge and when to rely on their own understanding to solve problems. Specifically, at each stage, we adopt GRPO as the base RL algorithm. For each query $x$, the attending physician generates a group of $G$ responses $\{y_{\text{final}}^{(1)},y_{\text{final}}^{(2)},...,y_{\text{final}}^{(G)}\}$, where each $y_{\text{final}}\sim\pi_{\theta_{\text{GP}}^{\text{attend}}}(y\mid x,y_{d})$. Subsequently, for each query with $G$ responses, GRPO computes the relative advantage of each response based on their rewards, which are determined by a format reward $R_{\text{format}}\in\{0,0.5\}$ and an accuracy reward $R_{\text{accuracy}}\in\{0,1\}$, avoiding training a complex critic model. The relative advantage is computed as $\small A_{i}=\frac{R_{i}-\text{mean}(\{R_{j}\}_{j=0}^{G})}{\text{std}(\{R_{j}\}_{j=0}^{G})}$. Then, as illustrated in Equation 1, GRPO uses the computed advantages to optimize the policy via the PPO-clip loss function, directly incorporating a KL penalty term to improve model performance. Notably, we apply different initial KL coefficients for tasks of varying difficulty levels, ensuring the model can gradually adapt to scenarios reflecting the performance of multiple specialist doctors.
Input: Task dataset $\mathcal{D}=\{x_{v}^{(i)},x_{t}^{(i)},y^{*(i)}\}_{i=1}^{N}$, policy model $\pi_{\theta}$, old policy $\pi_{\text{old}}$, group size $G$, responses from specialists $y_{d}^{(i)}$.
Output: $\pi_{\theta}$.
1 Initialize $\mathcal{D}_{\text{easy}}$, $\mathcal{D}_{\text{medium}}$, $\mathcal{D}_{\text{hard}}$ as empty sets
2 foreach $(x_{v},x_{t},y^{*})\in\mathcal{D}$ do
3       $\triangleright$ Use Specialists’ Accuracy to Categorize the Dataset by Task Difficulty
4       Calculate the accuracy of the specialist doctor $s\leftarrow\text{Acc}(y_{d},y^{*})$
5       if $s=1$ then
6             Put $\{(x_{v},x_{t}),y^{*}\}$ into $\mathcal{D}_{\text{easy}}$
7
8      if $0\textless s\textless 1$ then
9             Put $\{(x_{v},x_{t}),y^{*}\}$ into $\mathcal{D}_{\text{medium}}$
10
11      if $s=0$ then
12             Put $\{(x_{v},x_{t}),y^{*}\}$ into $\mathcal{D}_{\text{hard}}$
13
14foreach $(x_{v},x_{t},y^{*})\in\{\mathcal{D}_{\text{easy}},\mathcal{D}_{\text{medium}},\mathcal{D}_{\text{hard}}\}$ in batch do
15       $\triangleright$ Utilize the GRPO for Optimization at Each Stage
16       Sample $G$ rollouts $\{y_{\text{final}}^{(1)},y_{\text{final}}^{(2)},\cdots,y_{\text{final}}^{(G)}\}$ from $\pi_{\text{old}}$, where $y_{\text{final}}^{(g)}\leftarrow\pi_{\theta}(y\mid(x_{v},x_{t}),y_{d})$
17       foreach rollout $y_{\text{final}}$ do
18             Calculate the outcome reward $R(y_{\text{final}})=R_{\text{format}}(y_{\text{final}})+R_{\text{accuracy}}(y_{\text{final}})$
19      Compute the groupwise advantage $A_{i}\leftarrow\frac{R_{i}-\text{mean}(\{R_{j}\}_{j=0}^{G})}{\text{std}(\{R_{j}\}_{j=0}^{G})}$
20       Compute GRPO loss in Equation 1 and update $\pi_{\theta}$
Algorithm 1 Curriculum-Based Multi-Agent Reinforcement Learning (C-MARL)
4 Experiments
In this section, we evaluate the performance of MMedAgent-RL, aiming to answer the following questions: (1) Can MMedAgent-RL effectively improve model performance compared to other LVLMs and the Qwen2.5-VL-based baselines? (2) How does MMedAgent-RL perform on out-of-distribution datasets? (3) Does each proposed component contribute to performance gains? (4) What is the impact of choosing different models as specialist doctors? (5) Does MMedAgent-RL truly enhance the model’s capabilities across various specialist configurations?
4.1 Experimental Setup
Implementation Details. We use Qwen2.5-VL [3] (3B and 7B) as the base model. We design the prompt template using the format employed in MM-EUREKA [35], clearly specifying the required output structure, which includes using <think> and <answer> tags to separately contain the reasoning process and the final answer, with the two being separated. The detailed prompt is shown in Appendix. For training hyperparameters, the rollout batch size and training batch size are both set to 128, with 8 rollouts generated for each sample. The sampling temperature is set to 1.0 to encourage response diversity, and optimization is done with a learning rate of $1\times 10^{-6}$. Additionally, for the three stages of curriculum reinforcement learning, the KL divergence coefficients are set to $1\times 10^{-3}$, $4\times 10^{-3}$, and $1\times 10^{-2}$ respectively to stabilize training. For the number of specialists, we set $e=3$. For the framework, we adopt a multimodal RL framework based on OpenRLHF [17]. All training is conducted on 8 NVIDIA Tesla A100 80GB GPUs.
Baseline Methods. We compare MMedAgent-RL with methods under two different settings: 1) Single-agent setting: This includes a series of state-of-the-art LVLMs, encompassing both general LVLMs and domain-specific LVLMs. Specifically, we include comparisons of the LLaVA series [30], Yi-VL-34B [61], Qwen-VL [3], LLaVA-Med [24], MedFlamingo [36], RadFM [54], HuatuoGPT-Vision-7B [6] and GPT-4o [39]. 2) Multi-agent setting: This includes MedAgents [48], MDAgents [20], and SFT methods based on multi-agent feedback. To ensure a fair comparison, we use Qwen2.5-VL as the agent for decision making.
Data and Metrics. We train on the three medical VQA datasets, i.e., VQA-RAD [22], SLAKE [29], PathVQA [15]. Their test sets are considered the in-domain test sets. Additionally, following Chen et al. [6], we select the health and medicine subset of MMMU [62], and OmniMedVQA [18] as out-of-distribution datasets. All evaluation questions are multiple-choice, and accuracy is used as the evaluation metric.
4.2 Main Results
In this section, we conduct a comprehensive comparison on the medical VQA task involving six datasets and various LVLMs as well as baseline methods based on Qwen2.5-VL.
Comparison with Baselines in In-Distribution Datasets. Table 1 shows the performance of various models across four medical VQA benchmarks. General LVLMs like LLaVA-v1.6-34B and GPT-4o exhibit consistently strong performance, outperforming earlier medical-specific models such as Med-Flamingo and RadFM. Notably, GPT-4o achieves the highest average score (68.6%) among all single-agent models, demonstrating its powerful generalization capabilities even in specialized medical domains. Interestingly, the multi-agent collaboration strategy further boosts performance. MMedAgent-RL achieves the best overall average (72.7%), surpassing even the strongest single-agent models. This highlights the effectiveness of collaborative inference in leveraging the complementary strengths of different models. Among the collaborative approaches, using GPT-4o as the decision making agent also performs strongly across all benchmarks, further emphasizing the benefits of optimization of decision making agent in multi-agent systems.
Performance in Out-of-Distribution Datasets. We evaluate the performance of MMedAgent-RL across various out-of-distribution (OOD) datasets. The results are presented in Table 1, which demonstrates the generalization of our approach in adapting to different OOD scenarios. These two OOD datasets cover multiple body parts and involve various medical image modalities. Through reinforcement learning, MMedAgent-RL demonstrates significant superiority across multiple modalities, outperforming the base model by 15.3% and the SFT method by 25.1%. Moreover, it surpasses the performance of multi-agent collaboration methods that cannot optimize models, i.e., MedAgents and MDAgents, by 13% and 8%, highlighting the effectiveness of our approach in handling diverse and unseen data distributions.
Table 1: The results of the medical VQA benchmark. Here, MMMU denotes MMMU (Health & Medicine track). The best results are bold.
In-Domain Datasets Out-of-Distribution Datasets
Model VQA-RAD SLAKE PathVQA Avg. OmniMedVQA MMMU Avg.
GPT-4o 61.0 75.5 69.4 68.6 68.5 69.7 69.1
Med-Flamingo 45.4 43.5 54.7 47.9 30.7 28.3 29.5
RadFM 50.6 34.6 38.7 41.3 28.2 27.0 27.6
LLaVA-Med-7B 51.4 48.6 56.8 52.3 44.1 36.9 40.5
Qwen-VL-Chat 47.0 56.0 55.1 52.7 48.3 32.7 40.5
Yi-VL-34B 53.0 58.9 47.3 53.1 51.5 41.5 46.5
LLaVA-v1.6-7B 52.6 57.9 47.9 52.8 49.0 33.1 41.1
LLaVA-v1.6-13B 55.8 58.9 51.9 55.5 48.0 39.3 43.7
LLaVA-v1.6-34B 58.6 67.3 59.1 61.6 58.7 48.8 53.8
LLaVA-v1.5-LLaMA3-8B 54.2 59.4 54.1 55.9 44.6 38.2 41.4
HuatuoGPT-Vision-7B 63.0 77.2 58.7 66.3 74.6 51.0 62.8
Qwen2.5-VL-3B 61.0 62.7 57.6 60.4 60.1 54.5 57.3
Qwen2.5-VL-7B 61.8 64.7 60.5 62.3 60.8 56.6 58.7
Multi-Agent Collaboration
MedAgents 65.6 67.9 63.2 65.6 55.8 49.7 52.6
MDAgents 66.8 68.2 65.4 66.8 58.2 52.3 55.1
GPT-4o $\rightarrow$ Qwen2.5-VL-3B 61.8 60.3 49.3 57.1 52.0 46.2 49.6
GPT-4o $\rightarrow$ Qwen2.5-VL-3B+SFT 61.4 61.1 54.3 59.0 54.1 44.8 49.5
GPT-4o $\rightarrow$ Qwen2.5-VL-7B 62.5 63.9 53.2 59.9 56.4 50.7 53.6
GPT-4o $\rightarrow$ Qwen2.5-VL-7B+SFT 63.5 65.5 57.7 62.2 57.9 50.2 54.1
MMedAgent-RL (3B) 67.9 69.8 69.2 69.0 61.5 57.5 59.5
MMedAgent-RL (7B) 70.2 75.3 72.5 72.7 68.9 66.4 67.7
Figure 4: Results of different settings of specialist doctors.
Figure 5: Results under different levels of decision difficulty.
4.3 Analysis
In this section, we conduct a detailed performance analysis at each step and explore how model type, numbers of specialist doctors, and varying levels of decision difficulty affect the results, to better understand the performance gains achieved by MMedAgent-RL.
Ablation Studies. We conducted a series of ablation experiments to evaluate the impact of each component in MMedAgent-RL, as shown in Table 2. We can see that: (1) Reliable triage doctors are important. Accurately determining the department to which a specialist doctor belongs helps the model call upon knowledge from their corresponding field of expertise to answer questions, improving the accuracy of specialist doctors’ answers. A fine-tuned triage doctor significantly improves model performance compared to the original model, with an average performance increase of 3% across multiple datasets. (2) Based on this, the mechanism of specialist doctor consultation is introduced, further helping the decision-making agent fully utilize expert opinions, with an average performance increase of 4.5% across multiple datasets. (3) Most importantly, the addition of curriculum multi-agent reinforcement learning (C-MARL) enhanced the decision-making agent’s understanding of specialist doctors’ knowledge, achieving a significant performance improvement of 15.7%. This indicates that C-MARL can effectively solve the problem of overall misalignment between the model and external knwoledge. Specifically, each stage plays a corresponding role and can understand the specialist doctors’ knowledge according to the goals of different stages, achieving overall performance gains.
Table 2: Ablation results on ID and OOD datasets.
Model ID OOD
VQA-RAD SLAKE OmniMedVQA MMMU
MMedAgent-RL 67.9 69.8 61.5 57.5
w/o Triage 66.5 67.5 56.8 56.6
w/o Specialists 63.1 65.7 55.6 52.7
w/o C-MARL 61.8 60.3 52.0 46.1
+ Easy 64.5 66.4 56.8 55.7
    + Medium 66.2 68.4 57.4 56.6
     + Hard 67.9 69.8 58.5 57.5
Analysis of Specialist Doctors. We analyze the types and number of models playing the role of specialist doctors. Specifically, as shown in Figure 4, regarding the model types, the performance of the final decision-making agent is closely related to the performance of the specialist doctors. Therefore, we used a series of models that performed well on multiple datasets, such as GPT-4o [39], HuatuoGPT-Vision [6], and Qwen2.5-VL [3], as specialist doctors. Since the areas or tasks that each model excels in are not completely consistent, the specialist doctor played by GPT-4o ultimately performed the best. Its performance across various aspects was relatively balanced, enabling the final MMedAgent-RL to achieve the best performance.
Performance under Different Levels of Decision Difficulty. We show the performance of MMedAgent-RL on test data with different levels of decision difficulty in Figure 5. Since specialists cannot guarantee complete accuracy, this noise significantly impacts the model’s final decision-making. Therefore, through our C-MARL method, the model gradually learns to utilize specialist knowledge for decision-making during multi-stage training. The overall performance is 20% higher than the original model, especially on hard cases.
Case Study and Outlook on “Aha Moments". As shown in Figure 6, we demonstrate the model’s performance on multiple cases. We can clearly observe that MMedAgent-RL can provide accurate answers in the <answer> tags and also generate high-quality reasoning processes. We also found that this resembles the thought process of human doctors: it first defines the disease mentioned in the problem, carefully analyzes the images, and then determines whether it completely matches the disease definition. Moreover, it also analyzes the results provided by specialists. In our cases, it can accurately evaluate the specialists’ outputs and then reason out the correct answer. Please refer to more cases in Sec. E.4, with a greater variety of examples. Although there isn’t an “aha moment” seen in LLMs [13], this structured reasoning may inspire more human-like scientific AI systems in the future.
5 Related Work
Medical Vision-Language Models. The advancement of Vision-Language Models (VLMs) [30, 31, 64, 2, 8, 57] has catalyzed significant progress in medical applications [56, 6, 59, 58, 65], with large-scale models like LLaVA-Med [25], HuatuoGPT-Vision [6], and VILA-M3 [38] demonstrating profound results in medical diagnostics. Specialty-specific Med-VLMs have been deployed across various domains including XrayGPT [51] for radiology, PathChat [34] for pathology, and OphGLM [12] for ophthalmology. However, single models struggle to handle cross-domain expertise. Although multi-agent systems [27, 20, 48] have been proposed to combine diverse medical expertise, existing approaches typically use preset workflows that lack adaptive reasoning capabilities. Additionally, most Med-VLMs are developed through SFT on general VLMs using biomedical instruction data [6, 24, 32, 52], which is limited by the scarcity of high-quality reasoning examples and often results in models that struggle with complex diagnostic reasoning across specialties.
Reinforcement Learning for Multimodal Reasoning. To address the limitations of static multi-agent systems and overcome the constraints of supervised fine-tuning, Reinforcement Learning (RL) offers a promising alternative for optimizing medical reasoning. RL evolves from establishing foundational frameworks for learning from human preferences [47, 9, 66] to developing sophisticated approaches like RLHF for instruction following [40] and self-correction [21]. Recent advances with DeepSeek-R1 [13] demonstrate that LLMs can leverage RL to enhance reasoning capabilities in complex tasks without supervision, showing exceptional performance in mathematics and coding challenges [60]. This success has extended to multimodal reasoning [35, 45, 53, 7, 50, 63, 33, 28, 5, 11], including the biomedical domain [41, 46]. However, prior RL-based approaches for multimodal reasoning have primarily focused on optimizing a single model, leaving the potential of RL for enhancing multi-agent medical collaboration largely unexplored.
Figure 6: Several case analyses. In the model responses, blue text represents the process of reasoning about relevant medical knowledge based on the question, and red text represents the analysis of the answer provided by the specialists. The top-left and bottom-right examples come from in-domain datasets, VQA-RAD and PathVQA, while the other two come from out-of-domain datasets, OmniMedVQA and MMMU.
6 Conclusion
In this work, we present MMedAgent-RL, a novel RL-based framework for multi-agent collaboration in medical multimodal reasoning. Inspired by real clinical workflows, MMedAgent-RL utilizes a triage-and-referral system that combines open-source and proprietary LVLMs to simulate GP-specialist interactions. By introducing a curriculum reinforcement learning strategy, we enable the attending physician model to handle noisy or conflicting specialist inputs, learning when to rely on external knowledge versus internal reasoning. Extensive experiments demonstrate the effectiveness and generalizability of our approach across a diverse set of medical VQA datasets. Beyond its performance gains, MMedAgent-RL reveals a promising trajectory toward reasoning models that emulate human-like diagnostic thinking, bridging the gap between algorithmic efficiency and clinical realism. We hope this work encourages the broader adoption of structured, multi-agent, and RL-driven approaches for complex, high-stakes multimodal reasoning tasks.
Appendix
Appendix A Evaluated Models
We evaluate a series of state-of-the-art LVLMs and Multi-agent. The single-agent models include LLaVA [30], Yi-VL-34B [61], Qwen-VL [3], LLaVA-Med [24], MedFlamingo [36], RadFM [54] and GPT-4o [39]. The multi-agent frameworks include prior collaborative systems such as MedAgents [48] and MDAgents [20], as well as our proposed MMedAgent-RL framework that introduces reinforcement learning for adaptive multi-agent reasoning.
  • GPT-4o [39] is OpenAI’s latest multimodal large model that supports text, image, and audio inputs. It exhibits strong generalization across vision-language benchmarks and serves both as a single-agent baseline and as a specialist in our multi-agent settings.
  • Med-Flamingo [36] is a multimodal few-shot learner designed for the medical domain. Built upon OpenFlamingo, it is further pre-trained on biomedical image-text data from scientific literature. It enables few-shot medical visual question answering with minimal supervision.
  • RadFM [54] is a domain-specific foundation model tailored for radiology. It leverages large-scale radiology reports and domain-adaptive learning to improve zero-shot and few-shot performance on radiographic image understanding.
  • LLaVA-Med [24] extends LLaVA to the biomedical domain by fine-tuning with medical image-instruction pairs. It enhances medical reasoning and answer generation with limited supervision using domain-specific visual-textual alignments.
  • Qwen2.5-VL [3] is a versatile vision-language model developed by Alibaba. It supports high-quality OCR, multi-turn dialogue, and reasoning over complex multimodal inputs. It is used both as a strong single-agent baseline and as the foundation of agents in our proposed framework.
  • Yi-VL-34B [61] is a large-scale multimodal model from 01.AI. With 34 billion parameters, it offers high-capacity visual understanding and serves as a powerful open-source baseline across medical and general VQA tasks.
  • LLaVA [31, 30] are general-purpose vision-language models trained via visual instruction tuning. Evaluated in several sizes (7B, 13B, 34B), they serve as strong single-agent baselines in both in-domain and out-of-domain medical benchmarks.
Appendix B Evaluated Datasets
We employ three established medical vision-language datasets: VQA-RAD [22], SLAKE [29], and PathVQA [15]. Furthermore, to evaluate out-of-distribution performance, we incorporate the health and medicine subset of MMMU [62] along with OmniMedVQA [18].
  • VQA-RAD [22] is a manually constructed dataset containing 315 radiology images with 3,515 question-answer pairs. The images are distributed across head, chest, and abdomen regions, and include both open-ended and binary "yes/no" questions. Each image is associated with multiple clinically relevant questions generated by medical professionals. The dataset aims to facilitate the development of visual question answering systems for the medical domain.
  • SLAKE [29] is a semantically-labeled knowledge-enhanced dataset featuring 642 radiology images and over 14,000 question-answer pairs. It offers comprehensive annotations including masks for semantic segmentation and bounding boxes for object detection. SLAKE is bilingual (English and Chinese) and covers 12 diseases and 39 organs across various body parts. The dataset also incorporates a medical knowledge graph with 5,232 medical knowledge triplets to support knowledge-based reasoning.
  • PathVQA [15] is a pathology-focused dataset containing 32,799 open-ended questions from 4,998 pathology images. The dataset was created using a semi-automated pipeline to extract images and captions from pathology textbooks and generate question-answer pairs using natural language processing. PathVQA aims to support the development of AI systems capable of answering clinical questions about pathology images, with each question manually checked for correctness.
  • MMMU [62] (Health & Medicine subset) is part of the Massive Multi-discipline Multimodal Understanding benchmark. This subset contains approximately 1,752 test questions across five disciplines: Basic Medical Science, Clinical Medicine, Diagnostics and Laboratory Medicine, Pharmacy, and Public Health. The questions require college-level subject knowledge and deliberate reasoning, challenging models to perform expert-level perception and reasoning tasks.
  • OmniMedVQA [18] is a comprehensive medical VQA benchmark collected from 73 different medical datasets, featuring images across 12 different modalities and covering more than 20 distinct anatomical regions. All images are sourced from authentic medical scenarios, ensuring alignment with real-world applications. The benchmark provides a diverse evaluation platform for testing the capabilities of large vision-language models in medical image understanding and reasoning.
Appendix C Overview of the Baselines
We evaluate MMedAgent-RL against two main multi-agent baselines, MedAgents [48] and MDAgents [20]. These baselines represent state-of-the-art approaches in medical visual question answering.
  • MedAgents [48] establishes a zero-shot multi-agent collaboration framework that simulates real-world clinical workflows. The framework encompasses five critical steps: gathering domain experts, proposing individual analyses, summarizing analyses into a report, iterating over discussions until consensus is reached, and making a final decision. Different agents are assigned specific medical roles and collaborate to solve complex medical reasoning tasks. The framework relies on pre-trained large language models without additional fine-tuning, enabling natural dialogue-based interactions between agents. MedAgents demonstrates how specialized medical knowledge from different domains can be integrated through structured agent collaboration, providing a strong baseline for multi-agent medical reasoning.
  • MDAgents [20] advances multi-agent medical systems by introducing adaptive collaboration mechanisms. Unlike fixed collaboration patterns, MDAgents dynamically selects the most appropriate agent configuration and communication structure based on the specific medical task. This framework allows for more flexible interactions between general practitioners and specialist agents, optimizing the collaboration pattern for different types of medical queries. MDAgents incorporates mechanisms to resolve conflicts between different agent opinions and adapts the consultation workflow to match the complexity of the medical case, resulting in more robust decision-making across diverse medical scenarios.
Appendix D Experimental Setup
D.1 Data Statistics
The data used in this work is shown in Table 3 and involves five multimodal medical datasets: VQA-RAD, SLAKE, PathVQA, OmniMedVQA and MMMU (Health & Medicine track). Among them, three are used as in-domain datasets, with their training sets employed for model training. The remaining two are directly used as out-of-domain (OOD) testing datasets. The specific data volume for each dataset used at each stage of Curriculum-Based Multi-Agent Reinforcement Learning is detailed in Table 3.
Table 3: The results of the medical VQA benchmark. Here, MMMU denotes MMMU (Health & Medicine track) and the number of training and testing phase denotes the number of QA items for each phase.
Model All VQA-RAD SLAKE PathVQA OmniMedVQA MMMU
Train 12,176 940 1,681 9,555 / /
- Easy 8,321 498 1,284 6,539 / /
- Medium 1,409 160 114 1,135 / /
- Hard 2,626 281 275 2,070 / /
Test 15,153 251 416 3,362 11,124 150
D.2 Hyperparameter Settings
We use Qwen2.5-VL [3] as the base model. We design the prompt template using the format employed in MM-EUREKA [35], clearly specifying the required output structure, which includes using <think> and <answer> tags to separately contain the reasoning process and the final answer, with the two being separated. The detailed prompt is shown in Table 4. For training hyperparameters, the rollout batch size and training batch size are both set to 128, with 8 rollouts generated for each sample. The sampling temperature is set to 1.0 to encourage response diversity, and optimization is done with a learning rate of $1\times 10^{-6}$. Additionally, for the three stages of curriculum reinforcement learning, the KL divergence coefficients are set to $1\times 10^{-3}$, $4\times 10^{-3}$, and $1\times 10^{-2}$ respectively to stabilize training. For the number of specialists, we set $e=3$. For the baseline implementation, i.e., MedAgents [48] and MDAgents [20], we use Qwen2.5-VL as the agent for decision making to ensure a fair comparison between multi-agent baselines and MMedAgent-RL. For model training of 3B and 7B, we use DeepSpeed ZeRO2 and ZeRO3 respectively [43]. For the training framework, we adopt a multimodal RL framework based on OpenRLHF [17]. All training is conducted on 8 NVIDIA Tesla A100 80GB GPUs.
D.3 Prompt
The prompt for the fine-tuning of base model is shown in Table 4. In this prompt, we provide the question options, the input image, and $k$ expert answers. In the experiment, $k$ is set to 3. The model needs to first generate the reasoning process within the <think> tag, and then provide the final answer within the <answer> tag.
Table 4: Prompt template used for reinforcement learning fine-tuning.
Prompt Template: As the General Practitioner coordinating this case, review the specialist expertise to make a final decision. Answer from <Specialist>: <SpecialistAnswer>. <Question> Provide your final assessment. You need to first think about the reasoning process in the mind and then provide the user with the answer. The reasoning process and answer are enclosed within <think> </think> and <answer> </answer> tags, respectively, i.e., <think> reasoning process here </think><answer> answer here </answer>. The answer must be chosen from the given options.
Appendix E Additional Results
E.1 Ablation Analysis
E.1.1 Performance of Triage Doctor
The accuracy of the triage doctors is shown in Table 5. We used the data with definitive department labels as the evaluation target. From the results, we can observe that triage is not as challenging as answering complex medical diagnostic questions. Instead, department classification resembles a modality classification process. The original model already achieved an accuracy of over 80%, and after our fine-tuning, the model’s performance has reached a human-level standard on these datasets.
Table 5: The performance of triage doctor.
Model VQA-RAD SLAKE PathVQA
Qwen2.5-VL-3B 95.62 92.16 77.53
Qwen2.5-VL-7B 96.21 94.41 80.58
MMedAgent-RL (3B) 99.96 99.94 98.56
MMedAgent-RL (7B) 99.98 99.96 99.12
E.1.2 KL Divergence Coefficient
We conduct ablation experiments on the KL divergence coefficient at each stage, and the results are shown in Figure 7. We observe that in the first stage, as the KL divergence coefficient increases, the model’s performance tends to stabilize. This indicates that when training with simple data, where the specialist doctor’s answers are entirely correct, i.e., the model merely needs to learn to imitate. In this case, an additional KL divergence loss is required to constrain the policy model’s update steps, preventing it from changing too drastically; otherwise, it would become a model that simply copies the specialist’s answers. In the second stage, the optimal KL divergence coefficient is slightly larger than in the first stage, suggesting that the model needs some autonomy to explore its own direction. This becomes even more apparent in the third stage, where the optimal KL divergence coefficient is significantly higher. This is reasonable because, when the specialist doctor’s answers are entirely incorrect, it becomes very difficult for the model to generate an accurate response. If the KL divergence loss is too large in this stage, the model cannot explore effectively to find the correct answer. Therefore, in conclusion, different KL divergence coefficients need to be set for each stage of curriculum reinforcement learning to ensure optimal model performance.
Figure 7: Ablation of KL divergence coefficient.
E.2 Scaling Up the Model Parameters to 7B
Comparison with Baselines in In-Distribution Datasets. Table 6 also presents the performance of 7B-scale models across the same in-distribution medical VQA benchmarks. Overall, 7B models exhibit consistent improvements over their 3B counterparts. Notably, Qwen2.5-VL-7B outperforms Qwen2.5-VL-3B across all in-domain datasets, achieving an average score of 62.3% compared to 60.4%. This trend is further amplified in collaborative multi-agent setups. Specifically, MMedAgent-RL (7B) achieves a new state-of-the-art in in-domain settings with a 71.2% average, surpassing its 3B version by 3.3 percentage points and outperforming all other collaborative baselines. These results underscore the scalability of our collaborative learning strategy and reinforce the importance of model capacity in complex medical reasoning tasks.
Performance in Out-of-Distribution Datasets. In OOD settings, larger models also demonstrate improved generalization. As shown in Table 6, MMedAgent-RL (7B) achieves an average score of 60.7%, representing a 5.8% gain over the best-performing single-agent 7B model (Qwen2.5-VL-7B, 54.9%). Compared to its 3B counterpart, it improves by 4.3%, demonstrating that scaling not only benefits in-domain tasks but also enhances robustness to domain shifts. The performance margins over MedAgents and MDAgents are particularly notable—MMedAgent-RL (7B) outperforms these baselines by 8.2% and 5.6%, respectively—highlighting the crucial role of reinforcement learning in optimizing collaborative decision making in medical VQA under distributional shifts.
Table 6: The results of the medical VQA benchmark.
In-Domain Datasets Out-of-Distribution Datasets
Model VQA-RAD SLAKE PathVQA Avg. OmniMedVQA MMMU Avg.
GPT-4o 61.0 75.5 69.4 68.6 68.5 69.7 69.1
Med-Flamingo 45.4 43.5 54.7 47.9 30.7 28.3 29.5
RadFM 50.6 34.6 38.7 41.3 28.2 27.0 27.6
LLaVA-Med-7B 51.4 48.6 56.8 52.3 44.1 36.9 40.5
Qwen-VL-Chat 47.0 56.0 55.1 52.7 48.3 32.7 40.5
Yi-VL-34B 53.0 58.9 47.3 53.1 51.5 41.5 46.5
LLaVA-v1.6-7B 52.6 57.9 47.9 52.8 49.0 33.1 41.1
LLaVA-v1.6-13B 55.8 58.9 51.9 55.5 48.0 39.3 43.7
LLaVA-v1.6-34B 58.6 67.3 59.1 61.6 58.7 48.8 53.8
LLaVA-v1.5-LLaMA3-8B 54.2 59.4 54.1 55.9 44.6 38.2 41.4
HuatuoGPT-Vision-7B 63.0 77.2 58.7 66.3 74.6 51.0 62.8
Qwen2.5-VL-3B 61.0 62.7 57.6 60.4 60.1 54.5 57.3
Qwen2.5-VL-7B 61.8 64.7 60.5 62.3 60.8 56.6 58.7
Multi-Agent Collaboration
MedAgents 65.6 67.9 63.2 65.6 55.8 49.7 52.6
MDAgents 66.8 68.2 65.4 66.8 58.2 52.3 55.1
GPT-4o $\rightarrow$ Qwen2.5-VL-3B 61.8 60.3 49.3 57.1 52.0 46.2 49.6
GPT-4o $\rightarrow$ Qwen2.5-VL-3B+SFT 61.4 61.1 54.3 59.0 54.1 44.8 49.5
MMedAgent-RL (3B) 67.9 69.8 69.2 69.0 61.5 57.5 59.5
MMedAgent-RL (7B) 70.2 75.3 72.5 72.7 68.9 66.4 67.7
E.3 Detailed Results
Traditional Medical Imaging Evaluation.
Table 7 presents the accuracy of various models across five major medical imaging modalities in the OmniMedVQA benchmark. Our model (MMedAgent-RL) demonstrates strong generalization across all categories, achieving an average accuracy of 61.5%, significantly outperforming previous state-of-the-art models including LLaVA-v1.6-34B (58.7%) and Qwen2.5-VL-7B (60.8%).
Specifically, our method achieves 66.2% on microscopy images, indicating robust capability in processing fine-grained, high-resolution visual data typical of pathology slides. On MRI and CT modalities, MMedAgent-RLreaches 63.4% and 61.3%, respectively, outperforming strong baselines such as LLaVA-v1.6-34B and Yi-VL-34B by a wide margin. These results show that our model captures both structural and soft-tissue anatomical details effectively. In X-Ray, our method maintains competitive performance (65.0%) compared to high-performing models like HuatuoGPT-Vision-7B (80.3%), while achieving the highest accuracy on Ultrasound (51.7%) among all models, demonstrating robustness in handling noisy, low-contrast imaging modalities.
MMMU Health & Medicine Track.
In Table 8, our model again establishes new performance standards, achieving 57.5% overall accuracy on the MMMU Health & Medicine test set. Compared to existing large models such as Qwen2.5-VL-7B (56.6%) and HuatuoGPT-Vision-7B (51.0%), MMedAgent-RLdemonstrates clear advantages.
Notably, our model excels across all five sub-domains: scoring 64.7% in Basic Medical Science (BMS), 63.0% in Clinical Medicine (CM), 53.0% in Diagnostics and Laboratory Medicine (DLM), 56.3% in Pharmacy (P), and 50.5% in Public Health (PH). These results reflect a well-rounded capability across both foundational scientific understanding and applied clinical knowledge. In particular, performance in CM and P shows substantial improvement over single-agent baselines, suggesting that our model benefits from enhanced reasoning and domain transfer.
Taken together, these results confirm the effectiveness of our approach in both imaging-based and knowledge-based medical VQA settings, and highlight the potential of our method as a comprehensive solution for multimodal medical understanding.
Table 7: The accuracy of OmniMedVQA within different modalities (excluding FP, OCT, and Dermatology). CT: Computed Tomography, MRI: Magnetic Resonance Imaging, Mic: Microscopy Images, X-Ray: X-ray, US: Ultrasound.
Model CT MRI Mic X-Ray US Avg.
Med-Flamingo 34.6 27.5 28.1 30.1 33.2 30.7
RadFM 33.3 22.0 28.0 31.5 26.1 28.2
LLaVA-Med-7B 25.3 35.9 44.0 31.7 83.7 44.1
Qwen-VL-Chat 51.5 43.9 49.5 63.1 33.5 48.3
Yi-VL-34B 39.8 51.4 61.4 64.2 40.5 51.5
LLaVA-v1.6-7B 40.1 54.8 48.8 53.3 47.9 49.0
LLaVA-v1.6-13B 40.0 47.4 50.5 59.6 42.6 48.0
LLaVA-v1.6-34B 50.6 60.9 62.8 74.7 44.5 58.7
LLaVA-v1.5-LLaMA3-8B 33.0 53.8 48.4 56.6 31.2 44.6
HuatuoGPT-Vision-7B 65.6 72.7 77.5 80.3 76.7 74.6
Qwen2.5-VL-3B 60.5 64.2 66.6 68.9 40.4 60.1
Qwen2.5-VL-7B 62.0 68.3 70.7 68.9 34.3 60.8
Multi-Agent Collaboration
MedAgents 55.0 57.2 59.1 58.6 49.0 55.8
MDAgents 58.1 60.5 61.7 60.2 50.6 58.2
GPT-4o $\rightarrow$ Qwen2.5-VL-3B 57.3 50.2 60.6 65.0 26.7 52.0
GPT-4o $\rightarrow$ Qwen2.5-VL-3B+SFT 56.7 51.1 58.9 70.1 27.5 52.9
MMedAgent-RL (3B) 61.3 63.4 66.2 65.0 51.7 61.5
Table 8: Results on the test set for the MMMU Health & Medicine track. The Health & Medicine track is divided into five categories: BMS for Basic Medical Science, CM for Clinical Medicine, DLM for Diagnostics and Laboratory Medicine, P for Pharmacy, and PH for Public Health. Results are obtained by submitting to the official website.
Model BMS CM DLM P PH MMMU Health & Medicine
Med-Flamingo 29.6 28.1 24.8 25.3 31.2 28.3
RadFM 27.5 26.8 25.8 24.7 29.1 27.0
LLaVA-Med-7B 39.9 39.1 34.6 37.4 34.0 36.9
Qwen-VL-Chat 36.5 31.7 32.7 28.4 34.6 32.7
Yi-VL-34B 49.4 48.9 43.2 40.5 32.0 41.5
LLaVA-v1.6-7B 40.5 36.9 32.1 32.3 26.9 33.1
LLaVA-v1.6-13B 53.6 46.7 33.3 22.2 40.0 39.3
LLaVA-v1.6-34B 56.4 56.0 46.9 46.7 41.7 48.8
HuatuoGPT-Vision-7B 60.7 63.3 36.7 50.0 44.4 51.0
Qwen2.5-VL-3B 67.8 53.3 43.3 55.6 53.3 54.5
Qwen2.5-VL-7B 67.9 56.7 36.7 66.7 56.7 56.6
Multi-Agent Collaboration
MedAgents 60.0 55.0 45.0 50.0 52.9 52.6
MDAgents 62.5 57.5 50.0 55.0 50.5 55.1
GPT-4o $\rightarrow$ Qwen2.5-VL-3B 50.0 53.3 36.7 55.6 36.7 46.2
GPT-4o $\rightarrow$ Qwen2.5-VL-3B+SFT 50.0 50.0 26.7 55.6 43.3 44.8
MMedAgent-RL (3B) 64.7 63.0 53.0 56.3 50.5 57.5
E.4 More Cases
To further demonstrate the robustness and versatility of our proposed model in multimodal medical applications, we present additional representative cases in Figure 8, Figure 9, Figure 10, Figure 11, Figure 12 and Figure 13. These examples encompass various clinical scenarios and imaging modalities, providing a comprehensive illustration of the model’s ability to effectively integrate and interpret diverse types of medical data.
Figure 8: Case 1 (OmniMedVQA) & Case2 (PathVQA). In the model responses, blue text represents the process of reasoning about relevant medical knowledge based on the question, and red text represents the analysis of the answer provided by the specialists.
Figure 9: Case 3 (OmniMedVQA). In the model responses, blue text represents the process of reasoning about relevant medical knowledge based on the question, and red text represents the analysis of the answer provided by the specialists.
Figure 10: Case 4 (PathVQA). In the model responses, blue text represents the process of reasoning about relevant medical knowledge based on the question, and red text represents the analysis of the answer provided by the specialists.
Figure 11: Case 5 (VQA-RAD) & Case 6 (VQA-RAD). In the model responses, blue text represents the process of reasoning about relevant medical knowledge based on the question, and red text represents the analysis of the answer provided by the specialists.
Figure 12: Case 7 (PathVQA) & Case 8 (MMMU). In the model responses, blue text represents the process of reasoning about relevant medical knowledge based on the question, and red text represents the analysis of the answer provided by the specialists.
Figure 13: Case 9 (MMMU). In the model responses, blue text represents the process of reasoning about relevant medical knowledge based on the question, and red text represents the analysis of the answer provided by the specialists.